# Challenges and proposed solutions in modeling multimodal data: A systematic review

### v3

@@ Abstract
Multimodal data modeling has emerged as a powerful approach in clinical research, enabling the integration of diverse data types such as imaging, genomics, wearable sensors, and electronic health records. Despite its potential to improve diagnostic accuracy and support personalized care, modeling such heterogeneous data presents significant technical challenges. This systematic review synthesizes findings from 69 studies to identify common obstacles, including missing modalities, limited sample sizes, dimensionality imbalance, interpretability issues, and finding the optimal fusion techniques. We highlight recent methodological advances, such as transfer learning, generative models, attention mechanisms, and neural architecture search that offer promising solutions. By mapping current trends and innovations, this review provides a comprehensive overview of the field and offers practical insights to guide future research and development in multimodal modeling for medical applications.
@@
Multimodal data, Challenges, Fusion strategy, Modeling
@@
In healthcare, data comes in various modalities like electronic health records (EHR), imaging, genomics, and wearable devices. Integrating more than one of these modalities in a multimodal data analysis holds significant promise for improving our understanding and management of complex health conditions 6. One notable domain where multimodal data analysis can make a transformative impact is in the management of chronic diseases such as asthma and chronic obstructive pulmonary disease (COPD) 2. Here, medical imaging data, such as computerized tomography, provides insights into lung morphology and pathological changes. Wearable devices monitor physiological parameters and activity levels, offering real-time data on the patient’s disease trajectory. Environmental data, encompassing air quality and pollutants, helps identify external factors influencing respiratory conditions. Clinical records and genomic information contribute to personalized treatment strategies, while patient-reported outcomes capture subjective experiences, shaping a patient-centric approach. Combining several or even all of these modalities allows us to understand the disease and reveal interactions between different aspects.
@@
Modeling multimodal data is a challenging task that extends well beyond the analysis of individual modalities. Specifically, each modality may have different data types, formats, scales, levels of noise, and levels of interpretability. Moreover, modalities may exhibit varying degrees of completeness, sparsity, and dimensionality, making it challenging to integrate and analyze them together. In the context of modeling multimodal data, fusion techniques are of great importance. A fusion strategy involves combining and integrating the different modalities to create a unified representation that captures the complementary information from each modality. These techniques can be divided into four main categories: (i) early fusion, (ii) intermediate fusion, (iii) late fusion, and (iv) hybrid fusion. (i) Early fusion involves integrating information from different sources at the data level without learning any marginal representations or feature extraction, treating them as unimodal input. Early fusion offers simplicity by integrating modalities at the data level, however, it often assumes that all input modalities are directly compatible and equally informative. It can capture basic cross-modal relationships present in raw data, but it typically struggles to model more complex, high-level interactions between modalities. In particular, early fusion may be inadequate when modalities differ in dimensionality, sampling rates, or semantic content, such as combining image data with structured omics data, where aligning and weighting heterogeneous features is non-trivial. (ii) Intermediate fusion, in contrast, involves learning modality-specific representations to identify correlations within each modality before generating joint representations or making predictions. This approach offers flexibility in selecting different architectures for feature extraction from various modalities 3. By employing distinct architectures and layers for each modality, it can effectively handle heterogeneous data. Additionally, intermediate fusion is resilient to dimensionality imbalances 3 and can accommodate missing modalities without significantly affecting performance 3, 4. (iii) Late fusion, on the other hand, trains separate models for each modality and integrates their individual predictions at a later stage. This method is particularly well-suited for managing heterogeneous modalities and addressing dimensionality imbalances and missing data. However, it may be less effective for modalities with strong correlations, as it does not capture interactions between different data sources. In addition, in scenarios where specific information holds greater significance, using late fusion offers notable benefits. The model prioritizes the most important information it has, then adds more details from other sources to improve its understanding. This prioritization allows for a more nuanced integration of information, enhancing the model’s ability to make informed decisions and predictions. (iv) Hybrid fusion combines elements from the aforementioned methods, allowing for flexible processing across multiple stages. This approach is especially effective in scenarios with diverse data modalities, where early fusion can be used for closely related data types, and later fusion stages facilitate the integration of more distinct modalities. It provides the adaptability needed to meet the specific requirements of multimodal models. The integration of diverse data modalities into a multimodal model is visually represented in Figure 1. The figure displays the four primary fusion strategies (early, intermediate, late, and hybrid), while other less frequently used strategies, such as hierarchical fusion, probability-based fusion, and maximum correlation-based fusion, are not shown here for simplicity but are briefly discussed in Section 2.
@@
Selecting the optimal fusion technique is only one of several challenges in multimodal modeling 5. This systematic review aims to identify the broader set of challenges encountered in modeling multimodal data for medical applications and to synthesize the solutions proposed in the reviewed studies. By structuring the review around key challenges such as missing modalities, small datasets, dimensionality imbalance, interpretability, and fusion strategies, we provide a challenge-oriented perspective that highlights not only the methodological progress achieved but also the gaps that remain.
Several review articles have discussed the application of deep learning to analyze multimodal data. For instance, Mohsen et al. 6 explored integrating multiple data modalities with artificial intelligence models for a variety of applications, including combining medical imaging with data from electronic health records (EHRs). Similarly, Azam et al. 7 and Zhang et al. 8 solely focused on fusing different imaging modalities. Behrad and Abadeh 9 and Stahlschmidt et al. 3 primarily concentrated on integrating omics data with other modalities. Huang et al. 10 address the fusion of multimodal EHR data using traditional machine learning and deep learning methods. Pawłowski et al. 5 and Shaik et al. 12 focus on fusion techniques and feature extraction methods.
@@
The remainder of this paper is organized as follows: In Section 1, we present the methodology used for this systematic review, including the search strategy, inclusion and exclusion criteria, and data extraction process. Section 2 provides an overview of the challenges encountered in modeling multimodal data, including missing modalities, small sample sizes, dimensionality imbalance, interpretability, and identifying optimal fusion strategies. For each challenge, we summarize its impact on multimodal modeling and describe existing solutions reported in the reviewed studies. Section 3 discusses the key findings, highlights gaps in the literature, and reflects on limitations and future research directions. Finally, the paper concludes with Section 4, summarizing the main insights and contributions of this review.
1 Methods
1.1 Search strategy
To systematically identify relevant articles, a structured search strategy was employed (see Table 5). The search was conducted using two major databases, PubMed and manual search from Google Scholar. The time frame for the search was limited until October 2023 for PubMed (search date: 2023/10/09) and until end of 2023 for Google Scholar (search date: 2024/01/02). The strategy was designed to identify studies addressing challenges and methodologies in modeling multimodal data, with a focus on clinical applications. First, studies were required to explicitly mention “model fusion”, “data fusion,” or “multimodal” in their title, ensuring relevance to the topic of multimodal data modeling. To ensure comprehensive coverage, the Boolean operator OR was used within each keyword group to combine related terms. Next, the search included terms related to key research purposes, such as “disease prediction”, “diagnosis”, “clinical outcome prediction”, “modeling”, or “exacerbation”, to capture studies with clinical applications. Additionally, the search incorporated terms representing diverse data modalities, including “electronic health records”, “imaging data”, “wearable devices”, “longitudinal data”, “temporal data”, or “time-series data”, to ensure coverage of a wide range of multimodal datasets. To further refine the results, the term “challenges” was included to focus on studies discussing methodological or analytical obstacles related to multimodal data integration. Finally, the Boolean operator AND was used to combine these components. The final search query combined these elements to ensure a comprehensive yet focused selection of studies. In addition, a manual search was performed to identify any potentially relevant studies not captured in the database search. The reporting of this review follows the PRISMA 2020 guidelines page2021prisma.
1.2 Inclusion and exclusion criteria
@@ 2 Results
The yearly frequencies of the identified challenges are illustrated in Figure Figure 3. Specifically, since 2020, there was a strong increase in studies addressing challenges in modeling multimodal data (the number of studies in 2023 is lower as the search strategy was limited until October 2023). This upward trend indicates that research in this area will likely continue to expand, addressing new challenges as multimodal applications become more widespread.
@@
Abdelaziz et al. 52   Linear interpolation   Heterogeneity: Convolutional neural network
Aghili et al. 33   Gradient boosting
@@
Ahuja et al. 34   Guided topic priors by leveraging prior information from informative surrogate features   Scalability: time-efficient variational Bayesian inference implemented with parallelization in C++
@@
Alam et al. 70   Extracting high order interactions within and among data modalities: Introducing a semiparametric kernel machine approach, that captures pairwise and higher-order Hadamard product interactions across features from different modalities
@@
An et al. 53   Capturing the complex relationships and temporal dynamics among different modalities and account for the irregularities in patient visit timings: Proposing Time-Aware Multi-Type Data Fusion Representation Learning: Uses bidirectional long short-term memory and convolutional neural network networks to extract temporal and spatial patterns from different data types and applies a self-attention mechanism to combine and highlight key features across data types.
@@
Bahador et al. 54   High-dimensional data and intensive computational requirements: low-level fusion technique
@@
Bi et al. 61   Proposing the Clustering Evolutionary Random Forest to process high-dimensional data by clustering decision trees in a random forest and retaining only the most accurate ones.
@@
Brand et al. 62   Temporal dependencies: integrating temporal relationships using $\ell_{2,1}$-norm regularization to ensure that features are consistent over time, enabling the identification of longitudinal biomarkers
@@
Boehm et al. 35   Synthesizing complementary modalities of data with emerging multimodal artificial intelligence methods
Cai et al. 36   Vision Transformer
@@
Cai et al. 17   Frequency Domain Transformer model, which leverages multimodal data (histopathological images and genomic data)
Cai et al. 71   Multimodal generative adversarial networks
@@
Cremonesi et al. 18   Interoperability of Data for federated learning: Common Data Model and standardizing data representation and semantics across the different sources of data
@@
Dai et al. 19   Vision Transformer to fuse multi-modal features by capturing interactions and long-range dependencies   Loss of interaction between modalities: joint attention module based on the vision Transformer
@@
Dong et al. 55   Hypergraph laplacian regularization term that captures high-order relationships among subjects   Heterogeneity: high-order Laplacian regularized low-rank representation
Eshaghi et al. 76   Kernel-Based   Heterogeneity: Kernel-Based
@@
Golovanevsky et al. 37   Capturing interactions between modalities: Cross-modal bi-directional attention
Guan et al. 56   Knowledge distillation
@@
Hao et al. 68   Proposing a biologically interpretable pathway-based sparse deep neural network called Cox-PASN
Hatami et al. 38   Weighted Ensemble Fusion
@@
Hou et al. 20   Incorporating an online masked autoencoder into the hybrid graph convolutional network
@@
Jabeen et al. 15   Data augmentation, transfer learning   Features selection using the differential evolution and introducing probability-based fusion method
Karaman et al. 39   Re-weighted loss function
@@
Khan et al. 16   Features selection using the differential evolution and introducing maximum correlation-based fusion approach
@@
Krix et al. 21   Using the integrated gradients method to obtain feature importance and attention mechanisms to make predictions explainable   Capturing complex relationships: employing graph convolutional networks and graph attention networks
Kustowski et al. 57   Transfer learning
Lei et al. 74   Dimension reduction/Feature selection
@@
Li et al. 40   Identifying key features contributing to predictions through feature importance ranking methods
Li et al. 13   Limited depth in multimodal information mining: Hirarchical fusion
@@
Li et al. 22   Gradient-weighted class activation mapping method   Transfer learning   Focal loss function
@@
Li et al. 63   Heterogeneity: MixEHR, two-level probabilistic joint matrix factorization approach
Lim et al. 41   Transfer learning
@@
Liu et al. 64   Using multi-task deep learning models   Combining deep learning scores with radiologist-assessed imaging features and clinical factors in Cox models for transparent survival prediction   Heterogeneity: Multi-task learning
@@
Liu et al. 65   High Dimensionality: Multi-Task Feature Selection to jointly learn feature selection across modalities and remove irrelevant/redundant features
Liu et al. 77   Multiple kernel learning, Group Lasso regularization   Heterogeneity: Multiple kernel learning
@@
Madabhushi et al. 80   Based on non-linear dimensionality reduction methods   Heterogeneity: Kernel-Based
Martí-Juan et al. 23   Variational autoencoder
@@
Mao et al. 58   Focal loss function   Capturing intermodality interaction: Convolutional neural network, cross-modality interaction module
Ning et al. 24   Latent representation learning
Partin et al. 25   Data augmentation   Weighted loss function
Rabinovici-Cohen et al. 42   SHapley Additive exPlanations   Hetrogeneity: Stratification
@@
Rahaman et al. 26   Mid-fusion approach with an attention mechanism   Saliency analysis is applied to the trained model to identify the most important features from each modality
@@
Saad et al. 43   Deep convolutional neural network   Multi-task learning, Synthetic data generation   Heterogeneity: Multimodal feature fusion
@@
Shao et al. 27   Discovering significant features: Metric learning, Hilbert-Schmidt Independence Criteria
@@
Sheu et al. 44   Hybrid model using SHAP and deep learning based on dense layers-batch normalization
Shi et al. 28   Feature selection: Generative adversarial network
Shickel et al. 59   Transfer learning
@@
Sükei et al. 29   Hidden markov model and including information from unlabelled samples via transfer learning
Thung et al. 79   Low-rank matrix completion algorithm
Tiulpin et al. 69   Using Gradient-weighted Class Activation Mapping
Tiwari et al. 81   Semi-Supervised Multi-Kernel Graph Embedding
Tran and Liatsis 75   Combining one-class classification and Real Adaboost
@@
Vanguri et al. 45   Dynamic deep attention-based multiple-instance learning model that incorporates a masking mechanism   This masking mechanism dynamically adjusts attention weights to ignore missing modalities while effectively using the available data to support robust prediction
Vivar et al. 72   Matrix completion
@@
von Lühmann et al. 66   General linear model with temporally embedded Canonical Correlation Analysis
@@
Wang et al. 30   Modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure
@@
Wang et al. 46   Gradient-weighted class activation mapping, Shapley Additive exPlanations
@@
Wang et al. 47   knowledge-infused global-local data fusion model
Wang et al. 67   Knowledge distillation
@@
Xia et al. 31   Introducing an adaptive fusion layer that assigns weights to different modalities based on their importance
@@
Xin et al. 60   Interpretable Deep Multimodal Fusion Framework based on Deep Canonical Correlation Analysis
Xing et al. 32   Gradient-guided knowledge modulation scheme
Xu et al. 48   Shapley additive explanations   Variational autoencoder
@@
Xu et al. 49   Deep latent representation sequence learning   Recurrent neural network
Yoo et al. 50   Transfer learning, Gradient Boosting
Zhang et al. 14   Throughout fusion transformer
@@
Zhang et al. 78   High-order correlation among features: Feature Correlation Hypergraph
Ziegler et al. 51   Multimodal conditional generative adversarial networks
@@
Zille et al. 73   Extracting relationship between different modalities: Combining Lasso regression and sparse canonical correlation analysis in a collaborative regression framework to enforce cross-correlation in regression problems
@@
Classification (AUC / Acc) 71, 17, 80, 26 AUC $\approx$ 0.93 71, 17; Acc= 0.90–0.92 80, 26
Survival prediction (C-index / AUC) 13, 40, 24 27, 28 C-index= 0.72–0.78 24, 27; AUC= 0.71–0.77 28; C-index= 0.766 13, 40
Prediction 21, 25, 43, 45 AUC= 0.85 43; AUC= 0.80 45;
AUROC= 0.898 21; AUROC= 0.806 25
Joint (Class. + Regr.) 52, 62 RMSE= 10 52; RMSE= 16.2 ± 0.76 62
Representation learning 61 Acc= 0.90
Regression (RMSE) 73 Not reported
Detection 70 $p$-value=0.000001

### v4

@@ Abstract
Background: Multimodal data analysis has gained significant attention in healthcare research. Integrating diverse data sources, such as omics, wearable sensors, imaging, environmental, and electronic health records, offers a more comprehensive view of patient health and supports personalized care. However, modeling multimodal data poses several challenges that must be addressed to ensure accurate and reliable analysis. This systematic review aims to identify and analyze the challenges encountered in modeling multimodal data in clinical domains.
Methods: A systematic search was conducted in PubMed to identify original studies that modeled two or more data modalities for medical applications and explicitly addressed at least one modeling challenge. In addition, a manual search was performed in Google Scholar to capture further relevant studies. Studies were included if they addressed the use of multimodal data in clinical settings, focusing on data from diverse modalities. Papers that focused solely on multimodal imaging or included audio and video data were excluded during the title and abstract screening. A full-text review of 84 papers was then performed, leading to additional exclusions for studies that did not directly address modeling challenges specific to multimodal data in the clinical domain. Data extraction was performed to identify reported challenges along with their corresponding solutions or proposed approaches.
Results: This review included 69 studies (58 from the systematic search and 11 manually retrieved). Identified challenges in modeling multimodal data fall into five categories: missing modalities, dimensionality imbalance, fusion techniques, small datasets, and interpretability. Missing data affects model performance, addressed by multitask learning, matrix completion, and generative adversarial networks. Small datasets remain a critical challenge, with techniques such as data augmentation, transfer learning, and knowledge distillation showing promise in overcoming data scarcity. Interpretability, essential for clinical use, benefits from attention mechanisms and gradient-based attribution. Optimal fusion strategies, early, intermediate, and late fusion, are refined using neural architecture search.
Conclusions: This review outlines the main challenges in modeling multimodal data in clinical domains and summarizes emerging strategies to address them. By mapping current trends and innovations, it provides a comprehensive overview of the field and offers practical insights to guide future research and development in multimodal modeling for medical applications, thereby supporting improved data integration in clinical practice.
@@
Multimodal data, Challenges, Fusion strategy, Modeling, Clinical applications, Deep learning
Background
@@
In healthcare, data comes in various modalities like electronic health records (EHR), imaging, genomics, and wearable devices. Integrating more than one of these modalities in a multimodal data analysis holds significant promise for improving our understanding and management of complex health conditions Mohsen et al. [2022]. One notable domain where multimodal data analysis can make a transformative impact is in the management of chronic diseases such as asthma and chronic obstructive pulmonary disease (COPD) Acosta et al. [2022]. Here, medical imaging data, such as computerized tomography, provides insights into lung morphology and pathological changes. Wearable devices monitor physiological parameters and activity levels, offering real-time data on the patient’s disease trajectory. Environmental data, encompassing air quality and pollutants, helps identify external factors influencing respiratory conditions. Clinical records and genomic information contribute to personalized treatment strategies, while patient-reported outcomes capture subjective experiences, shaping a patient-centric approach. Combining several or even all of these modalities allows us to understand the disease and reveal interactions between different aspects.
@@
Modeling multimodal data is a challenging task that extends well beyond the analysis of individual modalities. Specifically, each modality may have different data types, formats, scales, levels of noise, and levels of interpretability. Moreover, modalities may exhibit varying degrees of completeness, sparsity, and dimensionality, making it challenging to integrate and analyze them together. In the context of modeling multimodal data, fusion techniques are of great importance. A fusion strategy involves combining and integrating the different modalities to create a unified representation that captures the complementary information from each modality. These techniques can be divided into four main categories: (i) early fusion, (ii) intermediate fusion, (iii) late fusion, and (iv) hybrid fusion. (i) Early fusion involves integrating information from different sources at the data level without learning any marginal representations or feature extraction, treating them as unimodal input. Early fusion offers simplicity by integrating modalities at the data level, however, it often assumes that all input modalities are directly compatible and equally informative. It can capture basic cross-modal relationships present in raw data, but it typically struggles to model more complex, high-level interactions between modalities. In particular, early fusion may be inadequate when modalities differ in dimensionality, sampling rates, or semantic content, such as combining image data with structured omics data, where aligning and weighting heterogeneous features is non-trivial. (ii) Intermediate fusion, in contrast, involves learning modality-specific representations to identify correlations within each modality before generating joint representations or making predictions. This approach offers flexibility in selecting different architectures for feature extraction from various modalities Stahlschmidt et al. [2022]. By employing distinct architectures and layers for each modality, it can effectively handle heterogeneous data. Additionally, intermediate fusion is resilient to dimensionality imbalances Stahlschmidt et al. [2022] and can accommodate missing modalities without significantly affecting performance Stahlschmidt et al. [2022], Lee and Van der Schaar [2021]. (iii) Late fusion, on the other hand, trains separate models for each modality and integrates their individual predictions at a later stage. This method is particularly well-suited for managing heterogeneous modalities and addressing dimensionality imbalances and missing data. However, it may be less effective for modalities with strong correlations, as it does not capture interactions between different data sources. In addition, in scenarios where specific information holds greater significance, using late fusion offers notable benefits. The model prioritizes the most important information it has, then adds more details from other sources to improve its understanding. This prioritization allows for a more nuanced integration of information, enhancing the model’s ability to make informed decisions and predictions. (iv) Hybrid fusion combines elements from the aforementioned methods, allowing for flexible processing across multiple stages. This approach is especially effective in scenarios with diverse data modalities, where early fusion can be used for closely related data types, and later fusion stages facilitate the integration of more distinct modalities. It provides the adaptability needed to meet the specific requirements of multimodal models. The integration of diverse data modalities into a multimodal model is visually represented in Figure 1. The figure displays the four primary fusion strategies (early, intermediate, late, and hybrid), while other less frequently used strategies, such as hierarchical fusion, probability-based fusion, and maximum correlation-based fusion, are not shown here for simplicity but are briefly discussed in Section Results.
@@
Selecting the optimal fusion technique is only one of several challenges in multimodal modeling Pawłowski et al. [2023]. This systematic review aims to identify the broader set of challenges encountered in modeling multimodal data for medical applications and to synthesize the solutions proposed in the reviewed studies. By structuring the review around key challenges such as missing modalities, small datasets, dimensionality imbalance, interpretability, and fusion strategies, we provide a challenge-oriented perspective that highlights not only the methodological progress achieved but also the gaps that remain.
Several review articles have discussed the application of deep learning to analyze multimodal data. For instance, Mohsen et al. [2022] explored integrating multiple data modalities with artificial intelligence models for a variety of applications, including combining medical imaging with data from electronic health records (EHRs). Similarly, Azam et al. [2022] and Zhang et al. [2020] solely focused on fusing different imaging modalities. Behrad and Abadeh [2022] and Stahlschmidt et al. [2022] primarily concentrated on integrating omics data with other modalities. Huang et al. [2020] address the fusion of multimodal EHR data using traditional machine learning and deep learning methods. Pawłowski et al. [2023] and Shaik et al. [2023] focus on fusion techniques and feature extraction methods.
@@
The remainder of this paper is organized as follows: In Section Methods, we present the methodology used for this systematic review, including the search strategy, inclusion and exclusion criteria, and data extraction process. Section Results provides an overview of the challenges encountered in modeling multimodal data, including missing modalities, small sample sizes, dimensionality imbalance, interpretability, and identifying optimal fusion strategies. For each challenge, we summarize its impact on multimodal modeling and describe existing solutions reported in the reviewed studies. Section Discussion discusses the key findings, highlights gaps in the literature, and reflects on limitations and future research directions. Finally, the paper concludes with Section Conclusion, summarizing the main insights and contributions of this review.
Methods
Search strategy
To systematically identify relevant articles, a structured search strategy was employed (see Table 5). The search was conducted using two major databases, PubMed and manual search from Google Scholar. The time frame for the search was limited until October 2023 for PubMed (search date: 2023/10/09) and until end of 2023 for Google Scholar (search date: 2024/01/02). The strategy was designed to identify studies addressing challenges and methodologies in modeling multimodal data, with a focus on clinical applications. First, studies were required to explicitly mention “model fusion”, “data fusion,” or “multimodal” in their title, ensuring relevance to the topic of multimodal data modeling. To ensure comprehensive coverage, the Boolean operator OR was used within each keyword group to combine related terms. Next, the search included terms related to key research purposes, such as “disease prediction”, “diagnosis”, “clinical outcome prediction”, “modeling”, or “exacerbation”, to capture studies with clinical applications. Additionally, the search incorporated terms representing diverse data modalities, including “electronic health records”, “imaging data”, “wearable devices”, “longitudinal data”, “temporal data”, or “time-series data”, to ensure coverage of a wide range of multimodal datasets. To further refine the results, the term “challenges” was included to focus on studies discussing methodological or analytical obstacles related to multimodal data integration. Finally, the Boolean operator AND was used to combine these components. The final search query combined these elements to ensure a comprehensive yet focused selection of studies. In addition, a manual search was performed to identify any potentially relevant studies not captured in the database search. The reporting of this review follows the PRISMA 2020 guidelines Page et al. [2021].
Inclusion and exclusion criteria
@@ Results
The yearly frequencies of the identified challenges are illustrated in Figure 3. Specifically, since 2020, there was a strong increase in studies addressing challenges in modeling multimodal data (the number of studies in 2023 is lower as the search strategy was limited until October 2023). This upward trend indicates that research in this area will likely continue to expand, addressing new challenges as multimodal applications become more widespread.
@@
Abdelaziz et al. [2021]   Linear interpolation   Heterogeneity: Convolutional neural network
Aghili et al. [2022]   Gradient boosting
@@
Ahuja et al. [2022]   Guided topic priors by leveraging prior information from informative surrogate features   Scalability: time-efficient variational Bayesian inference implemented with parallelization in C++
@@
Alam et al. [2018]   Extracting high order interactions within and among data modalities: Introducing a semiparametric kernel machine approach, that captures pairwise and higher-order Hadamard product interactions across features from different modalities
@@
An et al. [2021]   Capturing the complex relationships and temporal dynamics among different modalities and account for the irregularities in patient visit timings: Proposing Time-Aware Multi-Type Data Fusion Representation Learning: Uses bidirectional long short-term memory and convolutional neural network networks to extract temporal and spatial patterns from different data types and applies a self-attention mechanism to combine and highlight key features across data types.
@@
Bahador et al. [2021]   High-dimensional data and intensive computational requirements: low-level fusion technique
@@
Bi et al. [2020]   Proposing the Clustering Evolutionary Random Forest to process high-dimensional data by clustering decision trees in a random forest and retaining only the most accurate ones.
@@
Brand et al. [2020]   Temporal dependencies: integrating temporal relationships using $\ell_{2,1}$-norm regularization to ensure that features are consistent over time, enabling the identification of longitudinal biomarkers
@@
Boehm et al. [2022]   Synthesizing complementary modalities of data with emerging multimodal artificial intelligence methods
Cai et al. [2022]   Vision Transformer
@@
Cai et al. [2023]   Frequency Domain Transformer model, which leverages multimodal data (histopathological images and genomic data)
Cai et al. [2018]   Multimodal generative adversarial networks
@@
Cremonesi et al. [2023]   Interoperability of Data for federated learning: Common Data Model and standardizing data representation and semantics across the different sources of data
@@
Dai et al. [2023]   Vision Transformer to fuse multi-modal features by capturing interactions and long-range dependencies   Loss of interaction between modalities: joint attention module based on the vision Transformer
@@
Dong et al. [2021]   Hypergraph laplacian regularization term that captures high-order relationships among subjects   Heterogeneity: high-order Laplacian regularized low-rank representation
Eshaghi et al. [2015]   Kernel-Based   Heterogeneity: Kernel-Based
@@
Golovanevsky et al. [2022]   Capturing interactions between modalities: Cross-modal bi-directional attention
Guan et al. [2021]   Knowledge distillation
@@
Hao et al. [2019]   Proposing a biologically interpretable pathway-based sparse deep neural network called Cox-PASN
Hatami et al. [2022]   Weighted Ensemble Fusion
@@
Hou et al. [2023]   Incorporating an online masked autoencoder into the hybrid graph convolutional network
@@
Jabeen et al. [2022]   Data augmentation, transfer learning   Features selection using the differential evolution and introducing probability-based fusion method
Karaman et al. [2022]   Re-weighted loss function
@@
Khan et al. [2020]   Features selection using the differential evolution and introducing maximum correlation-based fusion approach
@@
Krix et al. [2023]   Using the integrated gradients method to obtain feature importance and attention mechanisms to make predictions explainable   Capturing complex relationships: employing graph convolutional networks and graph attention networks
Kustowski et al. [2021]   Transfer learning
Lei et al. [2016]   Dimension reduction/Feature selection
@@
Li et al. [2022]   Identifying key features contributing to predictions through feature importance ranking methods
Li et al. [2022]   Limited depth in multimodal information mining: Hirarchical fusion
@@
Li et al. [2023]   Gradient-weighted class activation mapping method   Transfer learning   Focal loss function
@@
Li et al. [2020]   Heterogeneity: MixEHR, two-level probabilistic joint matrix factorization approach
Lim et al. [2022]   Transfer learning
@@
Liu et al. [2020]   Using multi-task deep learning models   Combining deep learning scores with radiologist-assessed imaging features and clinical factors in Cox models for transparent survival prediction   Heterogeneity: Multi-task learning
@@
Liu et al. [2020]   High Dimensionality: Multi-Task Feature Selection to jointly learn feature selection across modalities and remove irrelevant/redundant features
Liu et al. [2014]   Multiple kernel learning, Group Lasso regularization   Heterogeneity: Multiple kernel learning
@@
Madabhushi et al. [2011]   Based on non-linear dimensionality reduction methods   Heterogeneity: Kernel-Based
Martí-Juan et al. [2023]   Variational autoencoder
@@
Mao et al. [2021]   Focal loss function   Capturing intermodality interaction: Convolutional neural network, cross-modality interaction module
Ning et al. [2023]   Latent representation learning
Partin et al. [2023]   Data augmentation   Weighted loss function
Rabinovici-Cohen et al. [2022]   SHapley Additive exPlanations   Hetrogeneity: Stratification
@@
Rahaman et al. [2023]   Mid-fusion approach with an attention mechanism   Saliency analysis is applied to the trained model to identify the most important features from each modality
@@
Saad et al. [2022]   Deep convolutional neural network   Multi-task learning, Synthetic data generation   Heterogeneity: Multimodal feature fusion
@@
Shao et al. [2023]   Discovering significant features: Metric learning, Hilbert-Schmidt Independence Criteria
@@
Sheu et al. [2022]   Hybrid model using SHAP and deep learning based on dense layers-batch normalization
Shi et al. [2023]   Feature selection: Generative adversarial network
Shickel et al. [2021]   Transfer learning
@@
Sükei et al. [2023]   Hidden markov model and including information from unlabelled samples via transfer learning
Thung et al. [2013]   Low-rank matrix completion algorithm
Tiulpin et al. [2019]   Using Gradient-weighted Class Activation Mapping
Tiwari et al. [2011]   Semi-Supervised Multi-Kernel Graph Embedding
Tran and Liatsis [2016]   Combining one-class classification and Real Adaboost
@@
Vanguri et al. [2022]   Dynamic deep attention-based multiple-instance learning model that incorporates a masking mechanism   This masking mechanism dynamically adjusts attention weights to ignore missing modalities while effectively using the available data to support robust prediction
Vivar et al. [2018]   Matrix completion
@@
von Lühmann et al. [2020]   General linear model with temporally embedded Canonical Correlation Analysis
@@
Wang et al. [2023]   Modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure
@@
Wang et al. [2022]   Gradient-weighted class activation mapping, Shapley Additive exPlanations
@@
Wang et al. [2022]   knowledge-infused global-local data fusion model
Wang et al. [2020]   Knowledge distillation
@@
Xia et al. [2023]   Introducing an adaptive fusion layer that assigns weights to different modalities based on their importance
@@
Xin et al. [2021]   Interpretable Deep Multimodal Fusion Framework based on Deep Canonical Correlation Analysis
Xing et al. [2023]   Gradient-guided knowledge modulation scheme
Xu et al. [2022]   Shapley additive explanations   Variational autoencoder
@@
Xu et al. [2022]   Deep latent representation sequence learning   Recurrent neural network
Yoo et al. [2022]   Transfer learning, Gradient Boosting
Zhang et al. [2023]   Throughout fusion transformer
@@
Zhang et al. [2014]   High-order correlation among features: Feature Correlation Hypergraph
Ziegler et al. [2022]   Multimodal conditional generative adversarial networks
@@
Zille et al. [2017]   Extracting relationship between different modalities: Combining Lasso regression and sparse canonical correlation analysis in a collaborative regression framework to enforce cross-correlation in regression problems
@@
Classification (AUC / Acc) Cai et al. [2018, 2023], Madabhushi et al. [2011], Rahaman et al. [2023] AUC $\approx$ 0.93 Cai et al. [2018, 2023]; Acc= 0.90–0.92 Madabhushi et al. [2011], Rahaman et al. [2023]
Survival prediction (C-index / AUC) Li et al. [2022, 2022], Ning et al. [2023], Shao et al. [2023], Shi et al. [2023] C-index= 0.72–0.78 Ning et al. [2023], Shao et al. [2023]; AUC= 0.71–0.77 Shi et al. [2023]; C-index= 0.766 Li et al. [2022, 2022]
Prediction Krix et al. [2023], Partin et al. [2023], Saad et al. [2022], Vanguri et al. [2022] AUC= 0.85 Saad et al. [2022]; AUC= 0.80 Vanguri et al. [2022];
AUROC= 0.898 Krix et al. [2023]; AUROC= 0.806 Partin et al. [2023]
Joint (Class. + Regr.) Abdelaziz et al. [2021], Brand et al. [2020] RMSE= 10 Abdelaziz et al. [2021]; RMSE= 16.2 ± 0.76 Brand et al. [2020]
Representation learning Bi et al. [2020] Acc= 0.90
Regression (RMSE) Zille et al. [2017] Not reported
Detection Alam et al. [2018] $p$-value=0.000001